IFF Master

An IFF Browser Version 1.2

July 27, 1994

Kay Drangmeister

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 Kay Drangmeister This document was created using TEX, ‘texinfo.tex’, and ‘makeinfo’.

1 General

IFF Master is a program that allows you to have a view inside the structure of IFF files. It is not yet another picture viewer or sound sample player, but it displays the internal entities (chunks) of a file. For example, the headers of pictures (ILBM) or sound samples (8SVX) are displayed in clear, so you can directly read the size and depth of the image or the sampling rate.

From time to time you will find an unknown file on your hard disk, let’s say a preference file. Since preference files are often stored in IFFormat, there is a chance you can directy examine the contens. There are some programs which write incorrect IFF files (e.g. the ILBM.CAMG chunk is a likely candidate), or which write additional information (e.g. copyright or author chunks). With IFF Master you can easily verify such cases.

In future versions (see Chapter 6 [History], page 6) there will be options to manipulate files, e.g. delete/edit/add texts inside ANNO chunks, or add new chunks like a DPI chunk for images.

Current features include:

• MUI application

• localized GUI (with AmigaOS 2.1 or later)

• Extensive chunk type library (currently 31 form types, 208 known chunks, 43 of these with comprehensive structure description)

• Chunk contents are presented alternatively as structure, text or hex dump

• Bit fields and enumeration types are displayed in clear

• Fixed point values are printed in decimal (e.g. 8SVX.VHDR.Volume)

• Callback hooks for special attributes, e.g. the Mode-ID inside the CAMG chunk is de-referenced

    (e.g. ‘PAL: Hires’).

IFF Master is distributed under the concept of freeware. Standard disclaimer applies to this program.

2 Installation

Nothing to it: just leave all files and the catalogs directory as they are in one directory, or alternatively copy the appropriate ‘iffmaster.catalog’ to into the system’s locale directory (i.e. ‘LOCALE:Catalogs/yourlanguage/’). By the way: if you mixed up catalog files for different languages, just use IFF Master to discover the catalog’s language…:)

System requirements are:

• AmigaOS 2.0, or AmigaOS 2.1 for localized GUI

• MUI version 2.0 (‘muimaster.library’ v7), See Section 5.1 [MUI], page 5.

If you wish to start IFF Master inside the user-startup, you should do this with ‘runback iffmaster iconified’. This will cause IFF Master to start as an AppIcon, so that Icons may be dragged onto this Icon.

3 Usage

The main window consists of 3 groups. The topmost one is the file specification group, the middle part ist the chunk list and below there are the action buttons.

In the current version there is no way to manipulate files, all concerning action buttons (e.g. ‘Add’, ‘Del’) are inactive.

3.1 File Specification

There are different ways to open a file. The most common one is by selecting ‘Project/Open...’ from the menu, or by clicking on the popup gadget beside the string gadget for getting a file requester. Alternatively, you can simply drag an icon on IFF Master’s main window (AppWindow). Furthermore, you can open the file which is currently in the clipboard by using the menu

item ‘Project/Open Clip’.

3.2 The Chunk List

The chunk list can show several types of information: the left column shows the chunk ID (e.g. BMHD), then there is the chunk type (e.g. ILBM), and at the right side there is the size of the chunk. Below the action buttons there is a cycle gadget labeled ‘Show’ that specifies the format of the chunk list. If it states ‘Description’ you will get a description of the chunk’s purpose (e.g. ‘Bitmap Header’), and ‘Contents’ will display a short summary of the chunk’s contents. You can advance the cycle gadget by pressing the SPACE key.

Because it is impossible to display the entire contents of a chunk in a single line you can get a comprehensive list by first selecting a chunk in the chunk list and then pressing ‘Info’ (you can as well double-click on the chunk). A new window containing the chunk’s contents will appear. You can display the contents either as structure, plain text or as a hex dump. Change the presentation by using the register above the contents list, or by pressing CURSOR LEFT and CURSOR RIGHT.

You need not close the contents window to display the contents of another chunk. Just click on that chunk in the chunk list or press CURSOR UP or CURSOR DOWN after having activated the chunk list via the TAB key. The new contents are displayed using the appropriate data type.

4 Problems?

Building hex dumps of very large chunks (e.g. ILBM.BODY) can take a long time. Therefore the number of bytes in a hex dump can be limited to a certain ammount. The default limit is 512 bytes, it can be adjusted or turned off in the Preferences (select ‘Settings/Set Prefs...’ from menu).

Some chunks (e.g. FORM) are container chunks with no contents, and therefore there is no presentation. Some other chunks (e.g. BODY) contain raw data, presentation of these chunks is limited to hex dump.

5 Credits

I like to thank:

Stefan Stuntz for MUI, See Section 5.1 [MUI], page 5.

Eric Totel for MUI-Builder

Kai Iske for Support and Betatesting

Thomas Reinhardt for his IFF files

5.1 MUI

                          This application uses

                        MUI - MagicUserInterface
                  (c) Copyright 1993 by Stefan Stuntz

MUI is a system to generate and maintain graphical user interfaces. With
the  aid  of  a  preferences program, the user of an application has the
ability to customize the outfit according to his personal taste.
MUI is distributed as shareware. To obtain a complete package containing
lots of examples and more information about registration please look for
a  file  called  "muiXXusr.lha"  (XX means the latest version number) on
your local bulletin boards or on public domain disks.
          If you want to register directly, feel free to send

                         DM 20.-  or  US$ 15.-
                                  to
                             Stefan Stuntz
                        Eduard-Spranger-Straße 7
                             80935 München
                                GERMANY

6 History

Version 1.0

First release Version 1.1

FIX version 1.0 was compiled with option 68030 - sorry! NEW many new chunk types NEW menu (moved ‘about’ and ‘prefs’ into the menu, gaining some space) NEW specifying file names at command line NEW loading files directly from clipboard NEW improved keyboard handling

Version 1.2 NEW fully localized GUI NEW chunk contents can be saved NEW AppWindow and AppIcon, command line option iconified NEW improved formatting capabilities for structure entries NEW sophisticated IFF parser, even brain-dead formats like EMOD.EMIC are handeled

appropriately NEW chunks now have nominal length, a message is displayed if a chunk is too short or

too long BUG Beta testers reported problems with de-iconify function when an icon was dropped

onto the AppIcon. I removed the said call, so you have to double-click the AppIcon after dropping an icon.

The program is not yet finished, several things like displaying huge hex dumps may be improved. There are many features left…

My to do list includes: hex dump, saveable preferences, save IFF, import and export chunk contents, maybe XPK support.

I would greatly appreciate new ideas and enhancement requests. If you discover any chunk unknown to IFF Master, please let me know (i.e. try to send me that (short) file or even a description of the chunk’s structure, if known). If you dare to translate the catalog to your language, please contact me, I will send you the ct file and the instructions.

7 Author

Kay Drangmeister Am Hubland 16, Apt. 1323 97074 Würzburg Germany T. ++49-931-702079 (night+day) kay@cip.informatik.uni-wuerzburg.de K.Drangmeister@insider.sub.de Chapter 8: Index 8

8 Index

8

8SVX.VHDR.Volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

A

action buttons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 altering files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 AppIcon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 AppWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

C

Chunk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 chunk ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 chunk list format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 chunk liste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 chunk type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Chunks without contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Container chunks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 contents of chunk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

D

Del . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 description. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

E

EMOD.EMIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

F

file manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 file specification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 format of chunk list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

G

General . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

H

hex dump. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

I

iconified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ILBM.CAMG.ModeID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

L

list format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

M

manipulation of files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 MUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

P

Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 presentation of chunk contents . . . . . . . . . . . . . . . . . . . . . . 3 Problems? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

S

Show . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

TAB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

U

Usage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 General. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1 File Specification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2 The Chunk List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

4 Problems?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5.1 MUI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

6 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 8 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8